Problem 9-16 Market Value Capital Structure Suppose the Schoof Company has this book value balance sheet: Current assets FixeThe notes payable are to banks, and the interest rate on this debt is 10%, the same as the rate on new bank loans. These bank loans are not used for seasonal financing but instead are part of the company's permanent capital structure. The long-term debt consists of 30,000 bonds, each with a par value of $1,000, an annual coupon interest rate of 8%, and a 15-year maturity. The going rate of interest on new long-term debt, rd, is 10%, and this is the present yield to maturity on the bonds. The common stock sells at a price of $50 per share. Calculate the firm's market value capital structure. Do not round intermediate calculations. Round your answers to two decimal places.
Answer #1

Short term debt value doesnot change =Current liabilities + Notes Payable =20,000,000+10,000,000 =30,000,000

Coupon of debt =8%*1000 =80
Number of years =15
YTM =10%
Price of Bond =PV of Coupons+PV of Par value =80*((1-(1+10%)^-15)/10%)+1000/(1+10%)^15 =847.8784
market value of debt =Price of debt* Number of bonds =847.87840987*30000=25436352.30

Market value of equity =Number of shares*Price of Shares =1000000*50 =50,000,000

Total capital =30,000,000+25436352.30+50,000,000=105436352.30

Short term capital% =30,000,000/105436352.30=28.45%
Long term capital% =25436352.30/105436352.30=24.12%
equity% =50,000,000/105436352.30 =47.42%
Total capital =30,000,000+25436352.30+50,000,000=105436352.30
Total capital =100%